Tier1 Operations Report 2018-06-11

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 11th June 2018

Review of Issues during the week 28th May to the 4th June 2018.
  • No incidents(major or minor), have been flagged during this reporting period.
Current operational status and issues
  • Continuing problems with small VOs and Castor (BDii) - Please be advised that in the light of the impact on a number of smaller VO's the removal of this service has caused, that Castor information (at least in the short term), is now being published in the BDii.
Resolved Castor Disk Server Issues
  • gdss738 - LHCb - d1t0. Currently back in production RO.
Ongoing Castor Disk Server Issues
  • gdss687 - LHCb - d1t0. Currently still in intervention.
Limits on concurrent batch system jobs.
  • CMS Multicore 550
Notable Changes made since the last meeting.
  • None.
Entries in GOC DB starting since the last report.
  • None
Declared in the GOC DB
  • None
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.

Listing by category:

  • Castor:
    • Update systems to use SL7 and configured by Quattor/Aquilon. (Tape servers done)
    • Move to generic Castor headnodes.
  • Networking
    • Extend the number of services on the production network with IPv6 dual stack. (Done for Perfsonar, FTS3, all squids and the CVMFS Stratum-1 servers).
  • Internal
    • DNS servers will be rolled out within the Tier1 network.
  • Infrastructure
    • Testing of power distribution boards in the R89 machine room is being scheduled for some time late July / Early August. The effect of this on our services is being discussed.

GGUS Tickets (Snapshot taken during morning of the meeting).

Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
135455 cms in progress less urgent 31/05/2018 04/06/2018 File Transfer Checksum verification at RAL EGI
135367 snoplus.snolab.ca in progress less urgent 28/05/2018 08/06/2018 Other Lost access to srm-snoplus.gridpp.rl.ac.uk EGI
135308 mice in progress top priority 24/05/2018 01/06/2018 Information System Can't send data to RAL Castor EGI
135293 ops on hold less urgent 23/05/2018 04/06/2018 Operations [Rod Dashboard] Issues detected at RAL-LCG2 EGI
134685 dteam in progress less urgent 23/04/2018 02/05/2018 Middleware please upgrade perfsonar host(s) at RAL-LCG2 to CentOS7 EGI
124876 ops on hold less urgent 07/11/2016 13/11/2017 Operations [Rod Dashboard] Issue detected : hr.srce.GridFTP-Transfer-ops@gridftp.echo.stfc.ac.uk EGI
GGUS Tickets Closed Last week
Request id Affected vo Status Priority Date of creation Last update Type of problem Subject Scope
135516 cms solved urgent 06/06/2018 08/06/2018 CMS_Facilities T1_UK_RAL WN-analysis and HammerCloud failures WLCG
135133 cms solved urgent 15/05/2018 05/06/2018 CMS_Data Transfers Likely corrupted File at T1_UK_RAL WLCG
134703 cms solved urgent 23/04/2018 07/06/2018 CMS_Data Transfers Transfer failing from RAL_Disk WLCG
133992 atlas verified less urgent 12/03/2018 04/06/2018 File Transfer RAL-LCG2-ECHO: No such file or directory EGI
127597 cms solved urgent 07/04/2017 07/06/2018 File Transfer Check networking and xrootd RAL-CERN performance EGI

Availability Report

Target Availability for each site is 97.0% Red <90% Orange <97%
Day Atlas Atlas-Echo CMS LHCB Alice OPS Comments
2018-06-04 100 100 97 100 100 0
2018-06-05 100 100 99 100 100 46.80851 Would hazzard a guess that BDii has just been put back!
2018-06-06 100 100 91 100 100 100
2018-06-07 100 100 99 100 100 100
2018-06-08 100 100 100 100 100 100
2018-06-09 100 100 99 100 100 100
2018-06-10 100 100 97 100 100 100
Hammercloud Test Report
Target Availability for each site is 97.0% Red <90% Orange <97%
Day Atlas HC CMS HC Comment
2018/06/04 95 39
2018/06/05 94 19
2018/06/06 89 84
2018/06/07 89 100
2018/06/08 95 89
2018/06/09 87 66
2018/06/10 89 66
2018/06/11 - -

Key: Atlas HC = Atlas HammerCloud (Queue RAL-LCG2_UCORE, Template 841); CMS HC = CMS HammerCloud

Notes from Meeting.
  • Notes from meeting of 13th June:
  • SKA:
    • It was noted that the SKA batch jobs are limited to 24 cores by the Dirac system (at Imperial) that they are using.
  • Dune/ProtoDune:
    • There is now a regular monday meeting for Dune/ProtoDune. Raja, as the Dune-Tier1 liaison person will attend regularly. Darren will attend in his place if Raja cannot be there.
    • Access to the Tier1 batch farm is being set-up. The Tier1's aim is to get this set-up for initial testing by next Monday's meeting. (18th June).
    • Alastair has set-up Dune access to Echo (line for access via WebDAV provided to Raja).
  • Euclid:
    • We are in contact with Euclid to get the public key for access to the Euclid CVMFS repository.

Documentation to Get VOs started accessing Tier1 Services: A start has been made on this. See: https://www.gridpp.ac.uk/wiki/RAL_Tier1_Echo